# EAS-SNN: End-to-End Adaptive Sampling and Representation for Event-based Detection with Recurrent Spiking Neural Networks

### v1

@@ 4 Experiments
The effectiveness of the proposed approach is initially substantiated in Sec. 4.1. Subsequently, in Sec. 4.2, we compare our comprehensive framework with state-of-the-art event-based methods across benchmark datasets. Additionally, Sec. 4.3 explores the adaptability, hyper-parameter sensitivity, and sampling diversity, while Sec. 4.4 delves into its energy efficiency.
Implementation Details. Our experiments are executed on four RTX 3090 GPUs. We utilize SpikingJelly [12] to enhance the training efficiency of the spike-based backbone. The ADAM optimizer is uniformly applied across all tests, with learning rates of 1e-3 for Gen1 and 1e-4 for N-Caltech 101. In alignment with practices from [19, 22], data augmentation techniques including random Zoom-in and horizontal flipping are employed for both datasets. For GEN-1, event sampling is capped at 200ms prior to the timestamp of labeling to manage data processing duration. Unlike the hard reset used in the sampling module, downstream networks implement a soft reset strategy. Rather than employing a learning rate scheduler, we leverage model weight smoothing during training through the use of an exponential moving average with a momentum of .
Benchmark Datasets. The Neuromorphic-Caltech 101 (N-Caltech 101) dataset [44], captured by an ATIS event camera [47] traversing 101 object categories via three saccades, comprises 8,246 event streams, each embodying 300 ms of saccades alongside bounding box annotations. We partition this dataset into an 8:2 split for training and validation. The Prophesee Gen1 dataset [10] presents another rigorous neuromorphic benchmark, offering over 39 hours of automotive recordings with more than 228k car and 28k pedestrian annotations for event-based detection in diverse driving scenarios. Adhering to the original evaluation protocol, prediction boxes with either a side length below 20 pixels or a diagonal shorter than 30 pixels are excluded after non-maximum suppression.
@@ 4.1 Ablation Study
In this section, we conduct a comparative analysis to ascertain the efficacy of individual components within the proposed sampling module as detailed in Table 1. In addition to comparing with the common event count approach, we introduce two novel embedding strategies leveraging SNNs and RSNNs, predicated on potential aggregation as detailed in Eq. 12. The results show that the integration of recurrent synaptic connections significantly enhances network performance on both datasets, with observed improvements of 6.7% on N-Caltech 101 and 4.9% on Gen1. Notably, employing the SAT and RPD modules in isolation with adaptive sampling sometimes results in a performance decline compared to RSNN (58.6% vs. 63.4% in mAP on N-Caltech 101). Conversely, when SAT and RPD are synergistically combined within the SNN framework, there is a significant performance improvement, yielding increases of 3.0% and 1.2% in mAP over RSNN on the N-Caltech 101 and Gen1 datasets, respectively. This demonstrates the critical importance of integrating these components to achieve optimal network performance.
| Method | RPD | SAT | N-Caltech 101 | Gen1 | ||
| mAP | mAP | mAP | mAP | |||
| Event Count | - | - | 0.587 | 0.369 | 0.611 | 0.321 |
| SNN | - | - | 0.567 | 0.367 | 0.670 | 0.366 |
| RSNN | - | - | 0.634 | 0.412 | 0.719 | 0.399 |
| ARSNN | 0.600 | 0.398 | 0.687 | 0.373 | ||
| ARSNN | 0.586 | 0.391 | 0.714 | 0.393 | ||
| ARSNN | 0.592 | 0.376 | 0.714 | 0.398 | ||
| ARSNN | 0.664 | 0.437 | 0.731 | 0.409 | ||
4.2 Performance Comparison
Next, we evaluate our proposed method against existing techniques on the Gen1 and N-Caltech 101 datasets. In addition to the proposed ARSNN in the embedding module, we adopt PLIF [14] neurons with learnable decay and tdBN [65] across the other spiking layers to improve the overall efficacy of SNNs. The models, sized at 8.92M and 25.3M parameters, mirror the configurations of YOLOX-S and YOLOX-M [17] respectively, achieved through scaling the network width. As evidenced in Table 2, our technique outperforms the state-of-the-art spike-based model [53] in performance metrics (0.354 vs. 0.310 mAP) while operating with a reduced parameter (8.92 M vs. 14.40 M) under three time steps. Even compared with ANNs that feature well-designed attention mechanisms [19] and recurrent architectures [32, 46], our strategy maintains competitiveness and notably marks the first instance of achieving over 40% COCO mAP on Gen1 through sparse spike-driven computations, thereby enhancing energy efficiency as detailed in Sec. 4.4. With regards to N-Caltech 101, to the best of our knowledge, our approach represents the first spike-based solution to attain competitive performance, setting a new benchmark in the field.
In object detection, modules are designed for specific roles: backbones extract features, Feature Pyramid Networks (FPN) fuse multi-scale features, and detection heads handle coordinate regression and object recognition. We examined the effects of sparse spike representation by integrating a non-spiking detection head and FPN into a hybrid architecture. As shown in Table 2, a spiking FPN within a compact 8.92M network outperforms its non-spiking counterpart, indicating that spike-based architectures might offer regularization advantages for event-driven detection. However, spiking detection heads within constrained time steps appear to reduce network performance.
| Method | Embedding | Backbone | Head | Temporal | #Params | Timestep | ||
@@ 4.2 Performance Comparison
| RVT [19] | HIST. | Transformer+RNN | YOLOX | Yes | 18.5M | 21 | - | 0.475 |
| EGO-12 [69] | EGO-12 | Transformer | YOLOv6 | No | - | - | - | 0.504 |
@@
| VC-MobileNet [8] | Voxel Cube | SNN | SSD | Yes | 12.64M | 5 | - | 0.174 |
@@
| EAS-SNN | ARSNN | SNN | YOLOX | Yes | 25.3M | 3 | 0.731 | 0.409 |
| YOLOX | Yes | 25.3M | 3 | 0.718 | 0.393 | |||
| YOLOX | Yes | 25.3M | 3 | 0.699 | 0.375 | |||
| EAS-SNN | ARSNN | SNN | YOLOX | Yes | 8.92M | 3 | 0.687 | 0.372 |
| YOLOX | Yes | 8.92M | 3 | 0.692 | 0.372 | |||
| YOLOX | Yes | 8.92M | 3 | 0.675 | 0.354 | |||
@@
| denotes the model uses a spiking backbone, a non-spiking feature pyramid network, and a non-spiking detection head. | ||||||||
| Method | Embedding | Backbone | Head | Temporal | #Params | Timestep | mAP mAP | |
@@
| EAS-SNN | ARSNN | SNN | YOLO-X | Yes | 25.3M | 3 | 0.664 | 0.437 |
| EAS-SNN | ARSNN | SNN | YOLO-X | Yes | 8.92M | 3 | 0.538 | 0.338 |
@@ 4.3 Experimental Analysis
In this section, our objective is to validate the efficacy of the proposed adaptive sampling technique within conventional dense neural networks. To this end, we configure to transmit only the initial aggregated potential to the subsequent networks. Fig. 4b displays the mAP training curves on the N-Caltech 101 dataset, comparing different methods. The combination of ARSNN with SAT and RPD demonstrates promising performance improvements, while ARSNN with SAT or RPD alone leads to declined performance, highlighting the synergistic benefits of our proposed modules. Additionally, introducing recurrent synaptic connections further enhances spike-based embedding performance. These results validate the adaptability of our sampling module for integration into dense neural networks.
@@
Impact of Early Aggregation: We investigate the influence of early aggregation, as defined in Eq. 4, by modulating the aggregation step on Gen1 using spiking YOLOX-S. Here, the first three slices sampling from -step inputs are aggregated and fed sequentially into a spike-based backbone. The parameter indicates the temporal resolution of inputs, where a higher corresponds to a finer granularity. As depicted in Fig. 4a, optimal performance is generally observed when remains consistent between the training and testing phases. Notably, the model trained with , sustains over 30% mAP, despite a fourfold reduction in granularity. Overall, models trained with demonstrate considerable robustness, delivering promising performance under all tested conditions.
@@
Distribution of Event Sampling: In this part, we employ the model trained under and examine the behavior of each neuron while using for inference. The sampling intervals of various spiking neurons are shown in the last column of Fig. 5, which demonstrates significant variability. The longest sample covers the full input, while the shortest sample only includes the input of one time step. Here, we disregard the neurons without spike firing, which typically make up due to the sparse event input. Additionally, compared to the event count representation in the first column, the image of the firing count in the second column offers a clearer estimation of the scene.
4.4 Energy Efficiency
@@
While ANNs execute a fixed number of multiplication-and-accumulation (MAC) operations determined by their architecture, SNNs on neuromorphic hardware utilize input-conditioned spikes for sparse accumulation (AC) operations, enhancing energy efficiency. To assess the energy efficiency of SNNs, we compute the average AC operation count across Gen1, as detailed in Table 4. The energy cost for 32-bit floating-point addition (AC) and multiply-and-add (MAC) are quantified at and per operation, respectively, following [23]. The results show a significant reduction in energy consumption ( for YOLOX-M and for YOLOX-S) with spike-driven computation. It is worth noting that, for ANNs, we omit the energy consumption of the embedding phase, offering a comparatively conservative estimate of power savings. The detailed firing rates and more results across different models can be found in Supplementary Materials.
@@ 5 Conclusion
We introduce the Adaptive Sampling technique with Recurrent Spiking Neural Networks (ARSNN), enhanced by Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) specifically to mitigate performance degradation in event-based sampling. Integrated with a spike-based architecture in the downstream neural network, our comprehensive framework, EAS-SNN, achieves superior performance relative to existing spike-based detection methods, delivering enhanced energy efficiency in comparison to ANNs. To the best of our knowledge, it is the first effort toward end-to-end optimization for event-based sampling and representation. Additionally, we demonstrate the versatility of our proposed sampling technique through its application to dense neural networks. Despite our exploration of the robustness across varying granularity of early aggregation, the ultimate goal is the development of a spike-based sampler that can precisely synchronize with the temporal granularity of raw event streams, measured in microseconds, without necessitating early aggregation. Our future work will focus on achieving this granularity, pending advancements in neuromorphic computing infrastructure.

### v2

@@ 4 Experiments
The effectiveness of the proposed approach is first validated in Sec. 4.1. In Sec. 4.2, we compare our framework with state-of-the-art event-based methods across benchmark datasets. Sec. 4.3 explores adaptability, hyper-parameter sensitivity, and sampling diversity, while Sec. 4.4 examines energy efficiency.
Implementation Details. The ADAM optimizer is applied with learning rates of 1e-3 for Gen1, 1e-4 for N-Caltech 101, and 5e-4 for 1Mpx. Consistent with [20, 24], data augmentation techniques, including random Zoom-in and horizontal flipping, are employed for all datasets. For Gen-1, event sampling is capped at 200ms prior to the timestamp of labeling. For 1Mpx, we preprocess the event stream following [20], while using a subsequence of length 4 for BPTT training. Unlike the hard reset used in the sampling module, downstream networks implement a soft reset strategy. Instead of employing a learning rate scheduler, we leverage model weight smoothing during training through the use of an exponential moving average with a momentum of . Additionally, we utilize SpikingJelly [13] to enhance the training efficiency of the spike-based backbone.
Benchmark Datasets. The Neuromorphic-Caltech 101 (N-Caltech 101) dataset [46], captured by an ATIS event camera [50], comprises 8,709 event streams, each containing 300 ms of saccades with bounding box annotations and object contours. We partition this dataset into an 8:2 split for training and validation. The Prophesee Gen1 dataset [10] offers over 39 hours of automotive recordings, annotated with more than 228k car and 28k pedestrian labels, facilitating event-based detection in diverse driving scenarios. The 1Mpx dataset [49] features high-resolution driving recordings (1280×720) across various scenarios, including over 24 million labeled bounding boxes of cars, pedestrians, and two-wheelers, annotated at 60Hz. Following the original evaluation protocol, we exclude prediction boxes with a side length below 10 (20) pixels or a diagonal shorter than 30 (60) pixels after non-maximum suppression for the Gen1 (1Mpx) dataset.
@@ 4.1 Ablation Study
In this section, we conduct a comparative analysis to ascertain the efficacy of individual components within the proposed sampling module, as detailed in Table 2. In addition to comparing with Event Count used in early aggregation, we introduce two novel embedding strategies leveraging SNNs and RSNNs, predicated on potential aggregation. The results show that the integration of recurrent synaptic connections significantly enhances network performance on both datasets, with observed improvements of 6.7% on N-Caltech 101 and 4.9% on Gen1. Notably, employing the SAT and RPD modules in isolation with adaptive sampling sometimes results in a performance decline compared to RSNN (58.6% vs. 63.4% in mAP50 on N-Caltech 101). Conversely, when SAT and RPD are synergistically combined within the SNN framework, there is a significant performance improvement, yielding increases of 3.0% and 1.2% in mAP50 over RSNN on the N-Caltech 101 and Gen1 datasets, respectively. This demonstrates the critical importance of integrating these components to achieve optimal network performance. Furthermore, Table 2 compares our approach with established event representations such as Voxel Grid [72], Time Surface [56], and Voxel Cube [8]. It shows that the proposed adaptive method outperforms these alternatives on both datasets. Moreover, the framework can incorporate various event representations at the early aggregation stage for better performance.
Method | RPD | SAT | N-Caltech 101 | Gen1 | |||
mAP50 | mAP50:95 | mAP50 | mAP50:95 | ||||
Event Count | - | - | 0.587 | 0.369 | 0.611 | 0.321 | |
SNN | - | - | 0.567 | 0.367 | 0.670 | 0.366 | |
RSNN | - | - | 0.634 | 0.412 | 0.719 | 0.399 | |
ARSNN | 0.600 | 0.398 | 0.687 | 0.373 | |||
ARSNN | 0.586 | 0.391 | 0.714 | 0.393 | |||
ARSNN | 0.592 | 0.376 | 0.714 | 0.398 | |||
ARSNN | 0.664 | 0.437 | 0.731 | 0.409 |
Method | N-Caltech 101 | Gen1 | |||
mAP50 | mAP50:95 | mAP50 | mAP50:95 | ||
Event Count | 0.587 | 0.369 | 0.611 | 0.321 | |
Voxel Grid | 0.598 | 0.391 | 0.515 | 0.247 | |
Time Surface | 0.497 | 0.288 | 0.654 | 0.354 | |
Voxel Cube | 0.589 | 0.401 | 0.651 | 0.340 | |
SNN | 0.567 | 0.367 | 0.670 | 0.366 | |
RSNN | 0.634 | 0.412 | 0.719 | 0.399 | |
ARSNN | 0.664 | 0.437 | 0.731 | 0.409 |
Method | Representation | Backbone | Head | Temporal | #Params | Timestep | ||
@@
RVT [20] | HIST. | Transformer+RNN | YOLOX | Yes | 18.5M | 21 | - | 0.472 |
ERGO-12 [75] | ERGO-12 | Transformer | YOLOv6 | No | - | - | - | 0.504 |
@@
VC-MobileNet [8] | Voxel Cube | SNN | SSD | Yes | 24.26M | 5 | - | 0.147 |
@@
EAS-SNN (M) | ARSNN | SNN | YOLOX⋄ | Yes | 25.3M | 3 | 0.731 | 0.409 |
YOLOX∗ | Yes | 25.3M | 3 | 0.718 | 0.393 | |||
YOLOX† | Yes | 25.3M | 3 | 0.699 | 0.375 | |||
EAS-SNN (S) | ARSNN | SNN | YOLOX⋄ | Yes | 8.92M | 3 | 0.687 | 0.372 |
YOLOX∗ | Yes | 8.92M | 3 | 0.692 | 0.372 | |||
YOLOX† | Yes | 8.92M | 3 | 0.675 | 0.354 | |||
@@
denotes the model uses a spiking backbone, a non-spiking feature pyramid network, and a non-spiking detection head. |
Method | Representation | Backbone | Head | Temporal | #Params | Timestep | mAP50 mAP50:95 | |
@@
EAS-SNN(M) | ARSNN | SNN | YOLO-X | Yes | 25.3M | 3 | 0.664 | 0.437 |
EAS-SNN(S) | ARSNN | SNN | YOLO-X | Yes | 8.92M | 3 | 0.538 | 0.338 |
4.2 Performance Comparison
We evaluate our proposed method against existing models on the Gen1, N-Caltech 101 and 1Mpx datasets. To improve the overall efficacy, we adopt PLIF [15] neurons with learnable decay and tdBN [70] in the downstream detector. The models, with sizes of 8.92M (SYOLOX-S) and 25.3M (SYOLOX-M) parameters, mirror the network architectures of YOLOX [18]. As shown in Table 3, our technique outperforms the state-of-the-art spike-based method [57] (0.354 vs. 0.310 mAP50:95) while using fewer parameters (8.92 M vs. 14.40 M) under three time steps. Even compared with ANNs that feature well-designed attention mechanisms [20] and recurrent architectures [34, 49], our method remains competitive and notably achieves over 40% mAP50:95 on Gen1 through sparse spike-driven operations, thereby enhancing energy efficiency as detailed in Sec. 4.4. Regarding N-Caltech 101 (Table 4), our approach improves the detection performance with spike-based representation. On the 1Mpx dataset, the largest event-based detection dataset, the proposed method achieves the first spike-based results, demonstrating remarkable performance using SNNs, as depicted in Table 5.
In object detection, each module has a specific role: backbones extract features, Feature Pyramid Networks (FPN) fuse multi-scale features, and detection heads handle coordinate regression and object recognition. We explored the impact of sparse spike representation by integrating a non-spiking detection head and FPN into a hybrid architecture. As shown in Table 3, a spiking FPN within a compact 8.92M network outperforms its non-spiking counterpart, indicating that spike-based architectures might offer regularization benefits for event-based detection. However, spiking detection heads within constrained time steps appear to reduce network performance.
Method | Representation | Backbone | Head | Temporal | #Params | Timestep | ||
RED [49] | Event Volum | CNN+RNN | SSD | Yes | 24.1M | - | - | 0.430 |
RVT [20] | HIST. | Transformer+RNN | YOLOX | Yes | 18.5M | 21 | - | 0.474 |
ERGO-12 [75] | ERGO-12 | Transformer | YOLOv6 | No | - | - | - | 0.406 |
ERGO-12 [75] | HIST. | Transformer | YOLOv6 | No | - | - | - | 0.327 |
E2VID [51] | Reconstruction | CNN | RetinaNet | No | - | - | - | 0.250 |
ASTMNet [34] | 1D TCNN | CNN+RNN | SSD | Yes | >100M | 3 | - | 0.483 |
Events+YOLOv3 [28] | HIST. | CNN | YOLOv3 | No | - | - | - | 0.346 |
EAS-SNN (M) | ARSNN | SNN | YOLOX | Yes | 25.3M | 3 | 0.651 | 0.362 |
@@ 4.3 Experimental Analysis
In this section, our objective is to validate the efficacy of the proposed adaptive sampling technique within conventional dense neural networks. Here, we only transmit the initial aggregated potential to subsequent networks. Fig. 4b displays the mAP50 training curves on the N-Caltech 101 dataset, comparing different methods. The combination of ARSNN with SAT and RPD demonstrates promising performance improvements, while ARSNN with SAT or RPD alone leads to declined performance, highlighting the synergistic benefits of our proposed modules. Additionally, recurrent synaptic connections further improve network performance.
@@
Impact of Early Aggregation: We investigate the influence of early aggregation, as defined in Eq. 4, by modulating the aggregation step on Gen1 using spiking YOLOX-S. The first three slices sampled from -step inputs are aggregated and fed sequentially into a spike-based backbone. The parameter indicates the temporal resolution of inputs, where higher corresponds to finer granularity. As depicted in Fig. 4a, optimal performance is generally observed when remains consistent between the training and testing phases. Notably, the model trained with , sustains over 30% mAP50:95, despite a fourfold increase in granularity. Overall, models trained with demonstrate considerable robustness, delivering promising performance under all tested conditions.
@@
Distribution of Event Sampling: In this part, we employ the model trained under and examine the behavior of each neuron while using for inference. The sampling intervals of various spiking neurons are shown in the last column of Fig. 5, which demonstrates significant variability. The longest sample covers the full input, while the shortest sample only includes the input of one time step. Here, we disregard the neurons without spike firing, which typically make up due to the sparse event input. Additionally, compared to the event count representation in the first column, the image of the firing count in the second column offers a clearer estimation of the scene. The qualitative results of adaptive sampling are further presented in Fig. 6. The top row shows event count image without adaptive sampling, under blurred and sparse textures. The bottom row demonstrates the improvement in information retention and noise reduction with adaptive sampling in the same contexts.
4.4 Energy Efficiency and Inference Speed
@@
While ANNs execute dense multiplication-and-accumulation (MAC) operations determined by their architecture, SNNs on neuromorphic hardware utilize input-conditioned spikes for sparse accumulation (AC) operations, thereby enhancing energy efficiency. To evaluate the energy efficiency of SNNs, we compute the average AC operations across Gen1, as detailed in Table 6. According to [25], we estimate the energy cost for 32-bit floating-point AC and MAC as and per operation, respectively. The results show a significant reduction in energy consumption with spike-driven computation, achieving a reduction for YOLOX-M and for YOLOX-S. It is worth noting that, for ANNs, we omit the energy consumption of the embedding phase, offering a comparatively conservative estimate of power savings. Detailed firing rates and additional results for different models are provided in Supplementary Materials.
Furthermore, we compared the frames per second (FPS) performance of our model with other state-of-the-art methods on the Gen1 dataset. Our model achieves competitive detection speeds of 54.35 FPS, compared to 59.88 FPS for RED [49], 28.09 FPS for ASTMNet [34], 106.38 FPS for RVT-T [20], and 59.52 FPS for GET [48]. This demonstrates the significant potential for higher efficiency on neuromorphic chips, which support sparse spike-driven operations.
@@ 5 Conclusion
We introduce an adaptive sampling method utilizing recurrent spiking neural networks (ARSNN), enhanced by Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) specifically to mitigate performance degradation in event-based sampling. Integrated with a spike-based architecture in the downstream neural network, our comprehensive framework, EAS-SNN, outperforms existing spike-based detection methods and offers superior energy efficiency compared to ANNs. To the best of our knowledge, it is the first effort toward end-to-end optimization for event-based sampling and representation. Additionally, we demonstrate the versatility of our proposed sampling technique through its application to dense neural networks. Despite exploring robustness across varying granularities of early aggregation, the ultimate goal is to develop a spike-based sampler that can precisely synchronize with the temporal granularity of raw event streams, measured in microseconds, without necessitating early aggregation. Our future work will focus on achieving this granularity, pending advancements in neuromorphic computing infrastructure.
Acknowledgements
This work was supported by National Natural Science Foundation of China under Grant 62236007, and the Key R&D Program of Zhejiang under Grant 2023C03001 and 2022C01048.